OpenStudy (anonymous):

area under the curve y=4-x^2, y=0. (Calc 2 question)

OpenStudy (agreene):

are you asking \[\int(4-x^2)dx\] or \[\int\limits_{0}^{0}(4-x^2)dx\]? the information you are providing seems a bit odd.

OpenStudy (anonymous):

im asking area under the curve y=4-x^2 bound by the line y=0. that would yeild an INT/limits from 1 to -1/(4-x^2)dx...I believe

OpenStudy (agreene):

Oh, I got yah know. So, start by finding the zeros of 4-x^2 (these will be your limits) -x^2=-4 x^2=4 x= 2;-2 \[\int\limits_{-2}^{2}4-x^2dx\] \[\int\limits_{-2}^{2}4dx-\int\limits_{-2}^{2}x^2dx\] \[4x-\frac{x^3}{3}\] take it to the limits: [4(2)-(2^3)/3]-[4(-2)-(-2^3)/3]=32/3